Thanks for your help! . . . . Sally made 24 cupcakes. She gave each of her 5 friends the same number of cupcakes, saving 4 for herself. Write an equation that could be used to find how many cupcakes each friend received.
We know the answer is 4 each . . .
(24-4)divided by 5= x
is this the right way to write that? Thank you!

Answer by mananth(16949)

Let each friend get x pancakes
there are 5 friends
she distributed a total of 5x pancakes.
she kept 4 with her.
5x+4 = 24( total number of pancakes)
5x= 24-4
5x=20
x= 20/5
x= 4